The average number of defects per 20 m2 area of a floor mat produced by a local company is two. A room has an area of 100 m2 and uses the same floor mat as described above.
i) Find the probability that there are at most two defects in a randomly selected 20 m2 of the floor mat.
ii) Find the probability that the floor mat contains at least three defects in the room.
1
Expert's answer
2019-11-13T11:32:19-0500

a) We need to use a Poisson distribution with "\lambda=2"


"P(k\leq 2 )=P(k=0)+P(k=1)+P(k=2)"

"P(k\leq 2 )=2^0\frac{e^{-2}}{0!}+2^1\frac{e^{-2}}{1!}+2^2\frac{e^{-2}}{2!}"

"P(k\leq 2 )=e^{-2}(1+2+2)=5e^{-2}=0.6767"

b) We need to use a Poisson distribution with "\lambda'=2\frac{100}{20}=2(5)=10"

"P(k\geq 3 )=1-P(k\leq 2 )"

"P(k\leq 2 )=10^0\frac{e^{-10}}{0!}+10^1\frac{e^{-10}}{1!}+10^2\frac{e^{-10}}{2!}"

"P(k\leq 2 )=e^{-10}(1+10+50)=61e^{-10}=0.0028"


"P(k\geq 3 )=1-0.0028=0.9972"